What is the total negative charge, in coulombs, of all the electrons in a small 1.05g sphere of carbon? One mole of C is 12.0g , and each atom contains 6 protons and 6 electrons.

Q = ?5.06
Answer #1

Part I

one mole to contain 6.02 x 10^23 carbon atoms, and it weighs 12.0g
This means that 1.00g of carbon contains 6.02 x 10^23/12.0 = 5.0 x 10^22 carbon atoms.
Knowing that each of the 5.0 x 10^22 carbon atoms contained in 1.00g has 6.00 electrons, then 1.00g of C has a total of 5.0 x 10^22 x 6.00 = 3 x 10^23 electrons.
Each of these electrons has a charge of -1.6 x 10^19 coulombs, so the total negative charge of1.00g of C = 3 x 10^23 electrons x -1.6 x 10^-19 coulombs/electron = -4.8 x 10^4 coulombs.

Answer: There's about -4.8 x 10^4 coulombs worth of negative charge in 1.00g of Carbon

Part II

First of all, what would remain on the other hand is protons and neutrons, only protons are charge and they have the same charge as the electrons but with a positive value so we can easily deduce that the total charge of what's remaining of the sphere is = +4.8 x 10^4 coulombs.
So let's pose ce to be the total charge of the electrons and cp the total charge of the protons, d the distance the separates them. And F the force between them.
ce = -4.8 x 10^4 coulombs
cp = +4.8 x 10^4 coulombs
d = 1.55m
So let's apply the formula:
F = k x I ce x cp I /d^2 K = 9 x 10^9
F = 9 x 10^9 x I -4.8 x 10^4 x +4.8 x 10^4 I / 1.50^2 = 9 x 10^9 x 4.8^2 x 10^8 / 1.50^2 = 9.216 x 10^18 N

Each of them will feel a force of about 9.2 x 10^18N

Part III

The force will be attractive because they have opposite charges.